HUBBLE SPACE TELESCOPE DAILY REPORT #957 PERIOD COVERED: 0800 EST Daily Status Report as of 236/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Complete WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were monitored in the OSS, and no anomalies were reported. The science data still remains on the science tape recorder. 1.2 Complete FOC/96 5030 (F/96 Filter Photometric Calibrations) The Faint Object Camera (f/96) was used to make observations of standard star HZ-4 to make calibrations of various photometric filters. A total of 11 observations were taken. They were monitored (when possible), received, and displayed. They all were reported to be normal. 1.3 Complete WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make two Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. Both observations were received, and both were reported to be okay. 1.4 Uplink Updated Ephemeris Tables, Per MOSES 2871 1.5 Complete PC 5014 (WF/PC 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make two standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled. One observations was reported to be partially saturated, and the other was normal. 1.6 Complete HSP/UV1/UV2/POL 5082 (X-Ray binaries Part 2 - Cycle 3 High) The High Speed Photometer was used to make six photometric and polarimetric observations of X-Ray binary star HD153919 (aka, 3U1700_37) to obtain more information about the structure and condition in the gas streams connecting the primary star to its neutron companion. All six observations were received and displayed in real-time. The target was reported to be well centered. 1.7 Complete HSP/UV2 4209 (Determination of the Masses and Densities of Pluto and Charon - cycle 3) The High Speed Photometer was used to make calibration observations of standard star BD+75D325 to determine the instruments long term sensitivity of the IDT detectors. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.8 Complete PC 4140 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor - Part 2) The WFPC was used to make internal flat calibrations images to check for new forms of contamination (i.e., Daises and measles), and to check for the possible return of quantum efficiency hysteresis. The observations completed as scheduled, and no problems were reported. 1.9 Complete HSP/POL 4887 (X-Ray Binaries - AUG - Cycle 3 High) The High Speed Photometer was used to make photometric and polarimetric observations of X-Ray binary star LMCX-3 to obtain more information about the structure and condition in the gas streams connecting the primary star to its neutron companion. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 8 Successful : 8 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Significant HSTARs: None 2.4 Operations Note: The Joint Integration Simulation #1 (JIS #1, Rendezvous to Solar Array Stow and EVA #1) was successfully initiated this morning (Tuesday Aug. 24) and will last through tomorrow (Wednesday Aug. 25). The test should last approximately 37 hours. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.2 Complete HSP/POL 4887 (X-Ray Binaries -AUG- Cycle 3 High) 3.3 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.4 Complete FOS 4108 (Supernova Remnants and Nucleosynthesis (FOS 30):Augmentation Cycle 2) 3.5 Complete HSP/POL/UV2 5082 (X-Ray Binaries Part 2 - cycle 3 High) 3.6 Complete HSP/POL/UV2 5070 (Optical and Ultraviolet Observations of Radio Pulsars - Cycle 2 High - Cont.) 3.7 Complete WFC 4848 (GTO Low Latitude Parallel Program - Cycle 3 High) 3.8 Complete FGS 5060 (Linking HIPPARCOS to the Extragalactic Reference Frame - Part 2 of 6) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-